The following set of points belong to a specific function: {(-3,0)(-2,4), (-1,0 ),(0,-6),(1,-8), (2,0),(3,24)} Based on the set of points answer the following questions:
a)What type of function does the set of points produce? Justify your answer.
b)Write an equation for this function based on the set of points that have been given.

Answer by stanbon(75887)

The following set of points belong to a specific function: {(-3,0)(-2,4), (-1,0 ),(0,-6),(1,-8), (2,0),(3,24)} Based on the set of points answer the following questions:
a)What type of function does the set of points produce? Justify your answer.
Passes thru the x-axis at (-3,0), (-1,0), (2,0)
Cubic
---------------------
b)Write an equation for this function based on the set of points that have been given.
f(x) = a(x+3)(x+1)(x-2)
f(-2) = 4 implies that 4 = a(1)(-1)(-4)
So 4a = 4
a = 1
------
Equation:: f(x) = (x+3)(x+1)(x-2)
------------------------------
Cheers,
Stan H.
-------------------